How do I solve this because I got stuck? The equation is: 14a=a^3-5a^2
Found 2 solutions by lynnlo, rothauserc:
Answer by lynnlo(4176)
Answer by rothauserc(4718)

we have
14a=a^3-5a^2
a^3-5a^2-14a=0
divide all terms by a on both sides of =, assumes a is not = 0
a^2-5a-14=0
(a-7)*(a+2) = 0
and a=7 or -2